OpenStudy (2yoututu11):

What is the approximate circumference of a semicircle with a radius of 12 centimeters? A. 18.8 centimeters B. 37.7 centimeters C. 75.4 centimeters D. 150.7 centimeters

OpenStudy (anonymous):

this is pi*r + 2r = 36 + 24 = 60

OpenStudy (anonymous):

oh - i have included the diameter - if u dont include that its B